# Online Optimization for Learning to Communicate over Time-Correlated Channels

### v4

@@ VII Conclusion
In this paper, we consider the problem of learning for communication over time-correlated channels via online optimization. To tackle this challenge, we employ the optimistic OMD framework to develop algorithms for designing communication systems. Furthermore, we provide theoretical guarantees for our methods by deriving sub-linear regret bounds on the expected error probability of learned communication schemes. Our theoretical findings confirm that devised algorithms utilize the distribution dependency of time-correlated channels to improve the performance of learned decoders and codebooks. To verify the effectiveness of our approaches, we conduct extensive simulation experiments and confirm the proposed methods’ superiority over other baselines, thus aligning with our theoretical discoveries.

### v5

@@ VI Simulation Results
(a) $\mu$ of GMD for decoder learning
(b) $\mu$ of LMD for decoder learning
(c) $\mu$ of GMD for codebook learning
(d) $\mu$ of LMD for codebook learning
Figure 5: For two considered tasks (online decoder and codebook learning), we explore the effect of different levels of channel correlation on the performance of our proposed optimistic OMD methods by setting different values of $\mu$.
[…]
In Figure 5 (a) and (b), we investigate the impact of $\mu$ on the performance of Euclidean-regularized optimistic OMD in decoder learning. We find that the average symbol error rate (SER) of optimistic OMD increases with $\mu$, indicating degraded performance as $\mu$ grows from $0.01$ to $0.07$. This is because a larger $\mu$ reduces the temporal correlation of channel gains $\mathbf{H}_{t}$, making them more statistically independent across rounds. Since optimistic OMD relies on leveraging channel correlation to enhance online learning, its performance deteriorates as $\mu$ increases. In Figure 5 (c) and (d), we explore the impact of $\mu$ on the performance of Log-barrier-regularized optimistic OMD in codebook learning. The empirical results in Figure 5 (c) and (d) are also consistent with our theoretical findings. In Figure 5 (d), although the SER with $\mu=0.07$ initially decreases faster, it ultimately remains higher than the SER with $\mu=0.01$.
(a) Baselines for Rayleigh channels
(b) Baselines for AWGN channels
Figure 6: For the online decoder learning task over Rayleigh fading channels, we compare our proposed Euclidean-regularized optimistic OMD algorithm with OGD and LS Estimation. For the online codebook learning task over AWGN channels, we compare our proposed Log-barrier-regularized optimistic OMD algorithm with EXP3 and Random selection.
To show the broad applicability of our proposed methods, we then conduct comparative experiments for online decoder and codebook learning in Rayleigh fading channels and AWGN channels, respectively. These two types of channels are commonly used to verify the performance of learning-based communication systems [3]. The channel simulation settings are consistent with those used in most previous studies on learning-based communication systems [3, 50]. The empirical results are provided in Figure 6.
From Figure 6 (a), we observe that our proposed optimistic OMD method surpasses both OGD and LS Estimation in learning decoders for Rayleigh fading channels. This improvement can be attributed to the additional update step in optimistic OMD, which enhances its performance compared to OGD. Moreover, optimistic OMD significantly outperforms LS Estimation by leveraging historical data more effectively through iterative updates. Figure 6 (b) further demonstrates that optimistic OMD outperforms baselines such as EXP3 and random selection in learning codebooks for AWGN channels.
@@ VII Conclusion
In this paper, we consider the problem of learning for communication over time-correlated channels via online optimization. To tackle this challenge, we employ the optimistic OMD framework to develop algorithms for designing communication systems. Furthermore, we provide theoretical guarantees for our methods by deriving sub-linear regret bounds on the expected error probability of learned communication schemes. Our theoretical findings confirm that devised algorithms utilize the distribution dependency of time-correlated channels to improve the performance of learned decoders and codebooks. To verify the effectiveness of our approaches, we conduct extensive simulation experiments and confirm the proposed methods’ superiority over other baselines, thus aligning with our theoretical discoveries. In future work, we will derive lower bounds for the problem of learning to communicate over time-correlated channels and develop novel algorithms capable of leveraging finer-grained channel information.